# Gradient Alignment Improves Test-Time Adaptation for Medical Image Segmentation

### v1

@@ Conclusion
In this paper, we proposed the GraTa, a TTA method to focus on optimizing the optimization direction and step-size during inference by gradient alignment and utilizing a dynamic learning rate, respectively. Specifically, we introduced an auxiliary gradient to the pseudo one to perform alignment to excavate the similarities between distinct gradients and further approximate the empirical gradient. The dynamic learning rate is constructed based on the cosine similarity between the pseudo and auxiliary gradients, encouraging pre-trained models to enhance learning as these gradients become closely aligned. Extensive experiment results on diverse scenarios demonstrated the superiority of our GraTa and the effectiveness of each special design. In our future work, we intend to explore the alignment of other auxiliary gradients to further improve the optimization procedure. We hope that our efforts can promote research on gradient optimization for test-time adaptation.

### v3

@@ Conclusion
In this paper, we proposed the GraTa, a TTA method to focus on optimizing the optimization direction and step-size during inference by gradient alignment and utilizing a dynamic learning rate, respectively. Specifically, we introduced an auxiliary gradient to the pseudo one to perform alignment to excavate the similarities between distinct gradients and further approximate the empirical gradient. The dynamic learning rate is constructed based on the cosine similarity between the pseudo and auxiliary gradients, encouraging pre-trained models to enhance learning as these gradients become closely aligned. Extensive experiment results on diverse scenarios demonstrated the superiority of our GraTa and the effectiveness of each special design. In our future work, we intend to explore the alignment of other auxiliary gradients to further enhance performance, as well as investigate other deployment scenarios, such as continual test-time adaptation (Chen et al. 2024). We hope that our efforts can promote research on gradient optimization for test-time adaptation.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Nos. 62171377, 92470101, 6240012686), in part by grants from National Key R&D Program of China (2022YFC2009903/2022YFC2009900), in part by the Ningbo Clinical Research Center for Medical Imaging under Grant 2021L003 (Open Project 2022LYKFZD06), in part by Shenzhen Science and Technology Program under Grants JCYJ20220530161616036, in part by the Fundamental Research Funds for the Central Universities (No. D5000230376), and in part by the Innovation Foundation for Doctor Dissertation of Northwestern Polytechnical University under Grant CX2024016.